f(x)=x^2-7x-8

What is the vertex, the line of symmetry, and the x and y intercept(s)? How do you find each one?
Thanks!

Answer by stanbon(75887)

f(x)=x^2-7x-8
What is the vertex, the line of symmetry, and the x and y intercept(s)? How do you find each one?
-------------------------
Vertex occurs at x = -b/(2a) = 7/2
f(7/2) = -20.25
---
Vertex::: (3.5,-20.25)
Line of symmetry: x = 3.5
--------
y-int = ?
Let x = 0, then y = -8
y-int = -8
-----------------
f(x)=x^2-7x-8
x-int = ?
Let y = 0
---
solve x^2-7x-8 = 0
(x-8)(x+1) = 0
x = 8 or x = -1
---
x-intercepts:
x = 8 and x = -1
======================
Cheers,
Stan H.
===========